OpenStudy (anonymous):

What is 5(2x+6)=-4(-5-2x)+3

OpenStudy (anonymous):

5(2x) + 5(6) = -4(-5)- 4(-2x) + 3 10x + 30 = +20 + 8x + 3 10x + 30 = +23 + 8x -10x -10x 30 = 23 -2x -23 -23 7 = -2x ------- -2 -2 -3.5 = x